OpenStudy (anonymous):

5(2x – 3) – (4x – 9) = 24

OpenStudy (anonymous):

answers cold be 14x – 19 = 24 6x – 6 = 24 6x – 24 = 24 6x – 24 = 24

OpenStudy (texaschic101):

5(2x - 3) - (4x - 9) = 24 -- distribute through the parenthesis 10x - 15 - 4x + 9 = 24 -- combine like terms 6x - 6 = 24 -- add 6 to both sides 6x = 24 + 6 6x = 30 -- divide by 6 x = 5 oops I did too much....answer is 6x - 6 = 24

OpenStudy (anonymous):

thanks man!

OpenStudy (texaschic101):

lol...its all good :)